# Probabilistic Neural Networks (PNNs) with t-Distributed Outputs: Adaptive Prediction Intervals Beyond Gaussian Assumptions

### v1

@@ 4 Experimental Results
We begin by presenting the results for the Concrete Compressive Strength data set, visualized in Fig. 6. Examining the coverage scores in Fig. 6(a), it is evident that none of the evaluated methods consistently achieve the target coverage level, indicating the inherent complexity and distributional challenges of this data set. This suggests that the Concrete Compressive Strength data deviates significantly from simple distributional assumptions, making accurate prediction interval estimation difficult. Focusing on coverage scores, TDistNNs demonstrate the highest coverage with 8 hidden neurons, and also achieve a superior maximum coverage score compared to GaussianNNs with 16 neurons, while GaussianNNs excel with 32 neurons. Notably, QuantileNNs consistently underperform, failing to reach the 90% coverage level across all configurations.
Figure 6: Boxplots illustrate (a) coverage score and (b) interval width for different architectures across 20 trials on the Concrete Compressive Strength data set.
@@
This issue is significantly exacerbated with GaussianNN. As Fig. 7(b) illustrates, some prediction intervals are entirely unreasonable, with upper bounds reaching values close to 1,000 and lower bounds dipping into negative territory—an impossibility given the strictly positive nature of the outputs. This observation aligns with our previous boxplot analysis, which demonstrated median average prediction interval widths well above 100, highlighting GaussianNN’s tendency to produce excessively wide and unrealistic intervals at the cost of reaching the target coverage level.
In contrast, Fig. 7(c) reveals that TDistNN’s constructed prediction intervals exhibit a stronger correlation with the true output values. For instance, when the true output is below 20, the majority of upper bounds remain below 40. Similarly, with a single exception, the upper bounds of TDistNN’s intervals consistently stay below 91.10—a marginal improvement over QuantileNN but a substantial advantage over GaussianNN. This analysis further underscores the critical importance of effectively modeling tail behavior in the predictive distribution. TDistNN’s ability to generate reasonable and adaptive prediction intervals demonstrates the practical benefits of robustly addressing heavy-tailed data, providing a more reliable and informative assessment of uncertainty compared to models relying on simplistic distributional assumptions.
Next, we evaluate the performance of these methods on the Energy Efficiency data set, where the output values range from 6.01 to 43.10. Fig. 8(a) presents the coverage scores, while Fig. 8(b) illustrates the corresponding interval widths across 20 independent trials. In terms of coverage, GaussianNNs achieve the highest scores, consistently exceeding the target coverage level of 90%. TDistNNs follow closely, with median coverage scores of 89.93, 88.31, and 88.31 for hidden layer sizes of 8, 16, and 32 neurons, respectively. Notably, TDistNN’s maximum coverage values (96.10, 96.75, and 97.40) indicate its ability to closely approach or even surpass the target coverage level. In contrast, QuantileNNs exhibit a progressive decline in coverage as the hidden layer width increases, with median coverage scores of 89.93, 87.98, and 84.74, suggesting potential instability in larger network configurations.
Figure 8: Boxplots illustrate (a) coverage score and (b) interval width for different architectures across 20 trials on the Energy Efficiency data set.
@@
Fig. 9 presents a comparative visualization of prediction intervals generated by the three methods across varying true output values. Notably, all three methods demonstrate strong performance when the true output values fall below 20. In this regime, however, the proposed TDistNN consistently produces the narrowest prediction intervals. This is particularly noteworthy given that the number of instances where the true target value falls outside the predicted interval is comparable across all three methods. This indicates that TDistNN achieves greater precision without sacrificing coverage.
@@
Due to the larger size of this data set, we increase the number of hidden units in the base neural networks to {8, 16, 32, 64, 128}, allowing for greater model capacity to capture complex patterns in the data. To ensure robustness and account for variability introduced by factors such as weight initialization, we repeat each experiment 20 times. This repeated evaluation helps mitigate the effects of randomness and provides a more reliable assessment of model performance. By scaling up both the data set size and network capacity, this experiment offers deeper insights into the behavior of the three studied methods, particularly their ability to generalize and maintain stable predictive performance in a larger-scale setting. The results are presented in Table 1, where we summarize the mean and standard deviation across 20 independent experiments.
Table 1: Comparison of coverage and interval width on the Student Performance Index data set. Results are presented as mean $\pm$ standard deviation.
QuantileNN GaussianNN TDistNN
neurons coverage width coverage width coverage width
8 $89.80\pm 0.36$ $10.06\pm 0.59$ $87.52\pm 5.47$ $194.74\pm 23.91$ $90.76\pm 0.13$ $6.80\pm 0.01$
16 $90.03\pm 0.37$ $8.39\pm 0.19$ $92.70\pm 2.10$ $195.71\pm 16.24$ $90.65\pm 0.12$ $6.77\pm 0.01$
32 $90.31\pm 0.31$ $7.50\pm 0.22$ $93.96\pm 1.41$ $186.58\pm 8.59$ $90.66\pm 0.18$ $6.75\pm 0.01$
64 $90.13\pm 0.39$ $7.08\pm 0.07$ $95.15\pm 0.89$ $169.68\pm 13.22$ $90.40\pm 0.20$ $6.73\pm 0.01$
128 $89.56\pm 0.53$ $6.81\pm 0.09$ $96.04\pm 0.91$ $150.28\pm 32.72$ $90.05\pm 0.22$ $6.69\pm 0.02$
Based on these results, we make the following observations:
@@
  • QuantileNNs achieve coverage scores close to or above the target level of 90% across all network configurations. Additionally, the variability in these coverage scores remains reasonably low, typically below 0.40, except for the largest network configuration (128 neurons), where slightly higher variability is observed. Regarding interval width, the prediction intervals produced by QuantileNNs are reasonable relative to the maximum true output value of 100. Notably, the mean interval width decreases substantially as the number of neurons increases, from 10.06 (with 8 neurons) to 6.81 (with 128 neurons), representing approximately a 47% reduction. This highlights that QuantileNN’s performance is noticeably influenced by the architecture of the underlying neural network.
  • Despite the larger sample size of this data set, GaussianNNs continue to exhibit difficulties in achieving a suitable balance between coverage and interval width. Although coverage scores occasionally surpass the target level of 90%, the corresponding interval widths remain excessively large—approximately 1.5 to 2 times the maximum true output value—indicating that the intervals are overly inflated. This issue arises primarily due to GaussianNN’s assumption of Gaussian-distributed outputs, which is often violated in practice. Additionally, we observe significantly greater variability in coverage scores for GaussianNN compared to QuantileNN, as evidenced by higher standard deviations, highlighting further instability in its predictions.
  • The proposed TDistNN consistently achieves coverage scores exceeding the target coverage level of 90% across all tested configurations. Additionally, the variability in these coverage scores, as reflected by their standard deviations, is notably lower compared to QuantileNN, indicating more stable and reliable performance. Furthermore, TDistNNs demonstrate the best overall trade-off between coverage and interval width on this data set, with prediction intervals exhibiting remarkable consistency—mean interval widths range narrowly from 6.69 to 6.80 with very small standard deviations. When normalized by the maximum true output value of 100, these interval widths remain below 7%, underscoring the model’s capability to produce precise and appropriately calibrated prediction intervals.
Consequently, TDistNNs provide a robust and effective approach to quantifying model uncertainty in practice, balancing accurate coverage with consistently narrow prediction intervals.

### v2

@@ 4.2 UCI Regression Benchmarks
We begin by presenting the results for the Concrete Compressive Strength data set, visualized in Fig. 5. Examining the coverage scores in Fig. 5(a), it is evident that none of the evaluated methods consistently achieve the target coverage level, indicating the inherent complexity and distributional challenges of this data set. This suggests that the Concrete Compressive Strength data deviates significantly from simple distributional assumptions, making accurate prediction interval estimation difficult. Focusing on coverage scores, TDistNNs demonstrate the highest coverage with 8 hidden neurons, and also achieve a superior maximum coverage score compared to GaussianNNs with 16 neurons, while GaussianNNs excel with 32 neurons. Notably, QuantileNNs consistently underperform, failing to reach the 90% coverage level across all configurations. As in the previous experiment, MC Dropout suffers from severe undercoverage in this case study, with median coverage ranging between 40% and 60%, well below the target coverage of 90%.
Figure 5: Boxplots illustrate (a) coverage score and (b) interval width for different architectures across 20 trials on the Concrete Compressive Strength data set. For each network size, the results are shown in the following left-to-right order: QuantileNN, GaussianNN, MC Dropout, and TDistNN.
@@
This issue is significantly exacerbated with GaussianNN. As Fig. 6(b) illustrates, some prediction intervals are entirely unreasonable, with upper bounds reaching values close to 1,000 and lower bounds dipping into negative territory—an impossibility given the strictly positive nature of the outputs. This observation aligns with our previous boxplot analysis, which demonstrated median average prediction interval widths well above 100, highlighting GaussianNN’s tendency to produce excessively wide and unrealistic intervals at the cost of reaching the target coverage level. Moreover, Fig. 6(c) illustrates that, despite producing more meaningful prediction intervals than GaussianNN, the resulting intervals exhibit clear undercoverage, with many failing to contain the ground-truth values.
In contrast, Fig. 6(d) reveals that TDistNN’s constructed prediction intervals exhibit a stronger correlation with the true output values. For instance, when the true output is below 20, the majority of upper bounds remain below 40. Similarly, with a single exception, the upper bounds of TDistNN’s intervals consistently stay below 91.10—a marginal improvement over QuantileNN but a substantial advantage over GaussianNN.
Next, we evaluate the performance of these methods on the Energy Efficiency data set, where the output values range from 6.01 to 43.10. Fig. 7(a) presents the coverage scores, while Fig. 7(b) illustrates the corresponding interval widths across 20 independent trials. With respect to coverage, GaussianNN attains the highest values, consistently surpassing the 90% target across all network sizes. In contrast, MC Dropout exhibits another limitation: although its coverage is relatively high for networks with 8 and 16 neurons, it drops sharply when the network size increases to 32 neurons, indicating that its performance is sensitive to the choice of network architecture.
TDistNNs provide reasonable coverage, with median coverage scores of 89.93%, 88.31%, and 88.31% for hidden layer sizes of 8, 16, and 32 neurons, respectively. Notably, TDistNN’s maximum coverage values (96.10%, 96.75%, and 97.40%) indicate its ability to closely approach or even surpass the target coverage level. In contrast, QuantileNNs exhibit a progressive decline in coverage as the hidden layer width increases, with median coverage scores of 89.93%, 87.98%, and 84.74%, suggesting potential instability in larger network configurations.
Figure 7: Boxplots illustrate (a) coverage score and (b) interval width for different architectures across 20 trials on the Energy Efficiency data set. For each network size, the results are shown in the following left-to-right order: QuantileNN, GaussianNN, MC Dropout, and TDistNN.
@@
Fig. 8 presents a comparative visualization of prediction intervals generated by the four methods across varying true output values. Notably, all four methods demonstrate strong performance when the true output values fall below 20. In this regime, however, the proposed TDistNN consistently produces the narrowest prediction intervals. This is particularly noteworthy given that the number of instances where the true target value falls outside the predicted interval is comparable across all four methods. This indicates that TDistNN achieves greater precision without sacrificing coverage.
@@ 4.3 Effect of Network Depth and Cross-Validation
Due to the larger size of this data set, we move beyond single-hidden-layer networks and consider deeper architectures with up to three hidden layers. We also vary the number of neurons in each hidden layer within the set $\{8,16,32\}$. Rather than relying on a single train-test split, we adopt a 5-fold cross-validation strategy and report the mean coverage and prediction interval width across folds for all four methods. In addition, we report the training time and prediction interval inference time as baseline measures of computational cost.
Because the loss function associated with the t-distributed output requires estimating the degrees-of-freedom parameter and computing additional elementwise operations (e.g., log, exp, softplus, and gamma functions), a modest computational overhead is expected, particularly for very small models (e.g., a single hidden layer with few neurons). By scaling up both the data set size and network capacity, this experiment provides deeper insight into the behavior of the studied methods.
Table 1: Comparison of coverage, prediction interval width, and computational time on the Student Performance Index data set. Results are reported as mean values across 5-fold cross-validation. Computational time is measured in seconds. For each row, the method achieving the lowest prediction interval width is highlighted.
QuantileNN GaussianNN MC Dropout TDistNN
Dep. Neurons Cov. Width Time Cov. Width Time Cov. Width Time Cov. Width Time
1 8 89.94 10.60 1.26 90.40 198.39 0.76 99.74 29.57 0.77 90.00 6.76 1.39
16 89.77 8.32 1.40 91.87 199.03 0.82 99.55 23.86 1.11 89.79 6.75 1.51
32 89.62 7.66 1.38 93.90 185.29 0.92 98.63 17.16 1.57 89.80 6.72 1.57
2 8 89.87 6.83 1.83 92.78 134.67 1.13 98.78 35.87 1.46 89.80 6.73 1.81
16 89.46 6.78 2.03 96.99 165.92 1.17 99.18 29.36 2.03 91.87 6.78 1.81
32 89.23 6.72 2.30 96.07 94.77 1.37 99.27 22.76 3.08 89.61 6.70 2.06
3 8 89.97 6.82 2.62 97.73 119.79 1.52 97.39 36.23 2.25 95.73 10.97 2.17
16 89.28 6.75 2.90 93.75 59.26 1.61 98.64 31.53 3.04 91.65 7.44 2.26
32 90.01 7.49 3.51 89.55 8.98 2.00 99.47 25.26 4.84 92.73 7.02 2.71
The results for a target coverage level of 90% are presented in Table 1. We then summarize key observations about the performance of the four models across different network depths and hidden-layer widths.
@@
  • QuantileNNs achieve consistent coverage scores ranging from 89.23% to 90.01% across all network configurations. Regarding interval width, the prediction intervals produced by QuantileNNs are reasonable relative to the maximum true output value of 100. Notably, when the network depth is fixed at 1, the mean interval width decreases as the number of neurons increases, from 10.60 (with 8 neurons) to 7.66 (with 32 neurons). In contrast, we observe less noticeable differences in interval widths when the network depth is increased to 2 or 3. Importantly, increasing the depth and the number of neurons per hidden layer does not necessarily always lead to smaller prediction intervals. For example, when the depth is 3 and the number of neurons is 32, the average predicted interval width increases to 7.49. This indicates that the performance of QuantileNNs is influenced by the architecture of the underlying neural network.
  • Despite the larger sample size of this dataset, GaussianNNs continue to exhibit difficulties in achieving a suitable balance between coverage and interval width. Although the coverage scores exceed the target level of 90%, the corresponding interval widths remain excessively large relative to the maximum true output value, indicating overly inflated prediction intervals. Interestingly, only one configuration with depth 3 and 32 neurons yields an average predicted interval width below 10, specifically 8.98. While this width is still larger than those produced by other models, such as QuantileNNs, it suggests that increasing the network depth or capacity can help mitigate the overestimation of variance in GaussianNNs. However, the downside is that the performance of GaussianNNs is overly sensitive to the network architecture.
  • Unlike previous experiments in which MC Dropout suffered from undercoverage, we observe a substantial increase in the mean coverage score across all configurations in this setting, with the minimum coverage reaching 97.39%, well above the target level of 90%. This improvement in coverage comes at the cost of significantly wider prediction intervals, which are approximately two to three times larger than those produced by QuantileNNs. One notable trend is that, for a fixed network depth, increasing the number of neurons tends to result in smaller prediction intervals, which is more desirable. Despite this trend, the minimum observed prediction interval width is 17.16, which remains uncompetitive compared to the other models under study. This also indicates that the performance of MC Dropout depends strongly on the network configuration and requires careful tuning.
  • The proposed TDistNN consistently achieves coverage scores close to or exceeding the target coverage level of 90% across all tested configurations. In contrast to the other models, the mean prediction interval widths are considerably more stable, remaining in the range of 6 to 7 for most configurations, with only one case reaching a maximum value of 10.97. Furthermore, for each configuration, corresponding to each row of Table 1, we highlight the model that achieves the smallest mean prediction interval width. In this comparison, TDistNN yields the smallest interval width in 7 out of 9 cases. Overall, these results indicate that the performance of TDistNN is less dependent on architectural configurations, which is a desirable property for practical applications where extensive hyperparameter tuning may be impractical. Across all configurations, TDistNN also achieves the minimum observed interval width of 6.70.
  • We also report the total computational time required to train each model across five runs corresponding to 5-fold cross-validation, together with the inference time. As expected, for all models, increasing the network size, either by increasing the depth or the number of neurons per hidden layer, leads to higher computational cost. When comparing different methods, GaussianNN is the fastest model, which can be attributed to its relatively simple loss function, the availability of a built-in loss implementation in PyTorch, and the fact that it does not require repeated training or sampling during inference, unlike QuantileNNs and MC Dropout. In terms of scalability as the network size increases, MC Dropout is the most computationally expensive approach due to the need to generate multiple stochastic forward passes at inference time. In our experiments, we use $T=100$ realizations, although this value can be adjusted to trade off accuracy and computational cost. While TDistNN is slightly more time-consuming than GaussianNN for a fixed configuration, a key advantage of TDistNN is that it achieves competitive coverage and prediction interval widths using smaller networks. For example, TDistNN requires 1.57 seconds to produce a mean interval width of 6.72 using a network with depth 1 and 32 neurons, whereas GaussianNN requires approximately 2 seconds to produce a wider prediction interval of 8.98 using a deeper network with depth 3 and 32 neurons.
Consequently, TDistNNs offer a robust and effective solution for practical uncertainty quantification, delivering accurate coverage and consistently narrow prediction intervals while maintaining stable performance across different network architectures.